Please, help me solve this problem:

The demand and supply functions for a newspaper company are, respectively, q = −10,000p + 3,100 and q = 2,000p + 700, where p is the price in dollars. At what price should the newspapers be sold so that there is neither a surplus nor a shortage of papers?

Answer by ankor@dixie-net.com(22740)

The demand and supply functions for a newspaper company are, respectively, q = −10,000p + 3,100 and q = 2,000p + 700, where p is the price in dollars.
At what price should the newspapers be sold so that there is neither a surplus nor a shortage of papers?
:
This means demand = supply, q = q, therefore
2000p + 700 = -10000p + 3100
2000p + 10000p = 3100 - 700
12000p = 2400
p = 2400/12000
p = $.20
Not that hard, right?